4. The most recent financial statements for Smith Corp. are shown below.


   Income Statement

Sales $16,200

Costs $10,400

Txbl Income    $5,800

Taxes (40%)    $2,320

Net Income    $3,480

   Balance Sheet

Current Assets $10,100 Debt    $16,400

Fixed Assets    $26,500 Equity $20,200

Total $36,600 Total    $36,600

Assets and costs are proportional to sales while debt and equity are not. The company maintains a constant 30% dividend payout ratio. What is the internal growth rate?

5. For Smith Corp. (above), what is the sustainable growth rate?
Answer #1

Retention ratio=1-payout ratio

=(1-0.3)=0.7

ROA=Net income/Total assets

(3480/36600)=0.0950819672

ROE=Net income/Equity

=(3480/20200)=0.172277228

internal growth rate=(ROA*Retention ratio)/[1-(ROA*Retention ratio)]
=(0.0950819672*0.7)/[1-(0.0950819672*0.7)]

=7.13%(Approx).

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.172277228*0.7)/[1-(0.172277228*0.7)]

=13.71%(Approx).